# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v2

@@ 7 Conclusion and Future Work
We introduce a general setting that embeds contextual information in a Bayesian setting. We showed that the regret can be lower bounded by $\Omega\left(\sqrt{TKI(\pi^{*};\mathcal{H}_{T})/\log K}\right)$. Combining this result with known upper bounds, this quantifies the relationship between a-priori external information and regret. Furthermore, in online settings these bounds quantify the relationship between online information accumulation and regret. We also utilized this trade-off in an MCQA task with LLMs, demonstrating that the aforementioned quantification can be easily used in an online setting.
@@
The general setting we introduced is relevant not only to the bandit setting, but also to general sequential decision problems. A central assumption in our analysis is that exogenous information can be gathered regardless of the process history. However, this assumption may be violated in general sequential decision process where the action itself may have ramifications on the quality, quantity and cost of the exogenous information.
@@
Our work only covers the Bayesian setting, and our definition of the measure of information relies on it. In other settings, such as adversarial learning [Neu and Olkhovskaya, 2020] or frequentist settings. Finally, and importantly, with the increasing prevalence of LLMs, and foundation models in general, building solid foundations as well as practical algorithms for using prior knowledge in sequential decision-making is an important research endeavor that may be built upon the foundations that are laid in this paper.
@@ Appendix A Table of Notations
$R_{M}(\pi)$ reward distribution according to decision $\pi$ on model $M$.
$\mu_{M}(\pi)$ mean reward according to decision $\pi$ on model $M$.
@@ Appendix B Useful Properties of Covering Numbers
Definition B.1 (Covering and packing numbers).
@@
  • A set $C\subset H$ is said to be a covering set of $H$ if for all $h\in H$ there is $c\in C$ such that $d(h,c)\leq\epsilon$.
@@
    $\mathcal{M}(d,\epsilon,H)=\max\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is a covering set of $H$}\right\}.$
@@
  • A set $C\subset H$ is said to be a packing set of $H$ if for all $h\in H$ there is $c\in C$ such that $d(h,c)\geq\epsilon$.
@@
    $\mathcal{N}(d,\epsilon,H)=\min\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is a packing set of $H$}\right\}.$
@@
  • We now define $B(\theta,\epsilon)=\{\theta^{\prime}\in H:d(\theta^{\prime},\theta)\leq\varepsilon\}$. $C$ is an $\epsilon$ local packing set of $B(\theta,\epsilon)$ if $\forall\theta^{\prime}\in B$, $\exists\theta^{\prime\prime}\in C$ such that $d(\theta^{\prime},\theta^{\prime\prime})\leq\frac{\epsilon}{2}$.
@@
    $\mathcal{M}^{\mathrm{loc}}(d,\epsilon,H)=\max\left\{m:\exists B(\theta,\varepsilon),\text{such that $C$ is an }\epsilon-\text{covering of $B$ and }\left|C\right|=m\right\}.$
@@
  • (Theorem 14.2 of Wu [2020]) For any subset of dimension $d$ that is contained in $\mathbb{B}_{1}^{d}$ the covering number $\mathcal{N}_{2}(\epsilon)$ obeys $\log\mathcal{N}_{2}(\epsilon)=\Theta(d\log(1/\epsilon))$. This includes the probability simplex with $d$ variables and $\mathbb{B}_{d}^{2}$.
  • (Lemma 6 of Gottlieb et al. [2017]) The covering of functions from $[0,1]^{d}$ to $[0,1]$ that are $L$-Lipschitz is given by $\log\mathcal{N}_{2}(\epsilon)=\Theta\left((L/\epsilon)^{d}\right)$.
@@
    $\mathcal{M}(\epsilon/2)\leq\mathcal{N}(\epsilon)\leq\mathcal{M}(\epsilon)$
@@ Appendix C Proofs for Section 3
$\mathcal{BR}^{*}(T)$ $=\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\mathbb{E}_{M\sim\nu}\left[\sum_{t=1}^{T}(\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t}))\right]$
$\geq\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\epsilon T\mathbb{P}_{\nu}\left(\sum_{t=1}^{T}(\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t}))\geq\epsilon T\right)$
$\geq\epsilon T\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\mathbb{P}_{\nu}\left(\forall t:\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t})\geq\epsilon\right)$
$\geq\epsilon T\sup_{\nu}\sup_{k\in\{1,\dots,K\}}\mathbb{P}\left(\forall t:\left|\mu_{M}(\phi_{k})-\mu_{M}(\pi_{t})\right|\geq\epsilon\right)$
$\geq\epsilon T\sup_{\nu}\sup_{k\in\{1,\dots,K\}}\mathbb{P}\left(\forall t:\rho(\phi_{k},\pi_{t})\geq\epsilon\right)$
$\geq\epsilon T\left(1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}\right)\,,$
where the last inequality follows from Fano’s inequality (Theorem 3.1). ∎
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}.$ (8)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq 2\epsilon^{2}T$. Furthermore, from Lemma 3.3, we have that $\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)\geq\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)$. Substituting this back into Equation 8,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+2\epsilon^{2}T}{\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)}.$
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}.$ (9)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq 2\epsilon^{2}T$. Furthermore, from Lemma 3.3, we have that $\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)\geq\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)$. Substituting this back into Equation 9,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+2\epsilon^{2}T}{\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)}.$
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (10)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)$. Substituting this back into Equation 10,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$
@@ C.3 Proofs for Additional Lower Bounds
Lemma C.1.
For any algorithm, there exists a Bayesian decision-making task with a finite decision such that the regret can be lower bounded by $\Omega(\sqrt{\left|\Pi\right|T})$.
@@
We focus on the set $\Delta(\Pi)$ and note that it is a parametric set. We also see that for all decisions $\phi,\psi\in\Pi$ we have that
$\mathbb{E}[\left|\mu_{M}(\phi)-\mu_{M}(\psi)\right|]\gtrsim\left\|\phi-\psi\right\|_{1}.$
Hence, in this scenario, we can take $\rho(x,y)=\left\|x-y\right\|_{1}$. Furthermore, using Lemma B.1 we have that
$\log\mathcal{M}_{1}\left(\epsilon/2,\Delta(\Pi)\right)-\log\mathcal{M}_{1}\left(\epsilon,\Delta(\Pi)\right)=\Theta\left(\left|\Pi\right|\right).$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\frac{1+2\epsilon^{2}T}{\left|\Pi\right|}.$
Now we fix $\epsilon$ such that $\frac{2\epsilon^{2}T}{\left|\Pi\right|}=\frac{1}{2}$. More specifically, we set $\epsilon=\frac{1}{4}\sqrt{\frac{\left|\Pi\right|}{T}}$ which yields
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim\frac{1}{2}\sqrt{\left|\Pi\right|T}$
which concludes our proof. ∎
The results for multi-armed bandits, contextual bandits, and tabular MDPs are shown by substituting $\left|\Pi\right|=A,CA$ and $SAH$ respectively.
Lemma C.2.
For any algorithm, the exists a Bayesian decision-making task with $\Pi\subseteq\mathbb{B}_{2}^{d}$, where the reward means are a linear function of the decisions, such that the regret can be lower bounded by $\Omega(\sqrt{dT})$.
@@
Similarly to Lemma C.1, we have $\rho(x,y)=\left\|x-y\right\|_{1}$. We focus on the set $\mathbb{B}_{2}^{d}$ and note that it is a parametric set. Furthermore, using Lemma B.1 we have that
$\log\mathcal{M}_{1}\left(\epsilon/2,\mathbb{B}_{2}^{d}\right)-\log\mathcal{M}_{1}\left(\epsilon,\mathbb{B}_{2}^{d}\right)=\Theta\left(d\right).$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\frac{1+2\epsilon^{2}T}{d}.$
Now we fix $\epsilon$ such that $\frac{2\epsilon^{2}T}{d}=\frac{1}{2}$. More specifically, we set $\epsilon=\frac{1}{4}\sqrt{\frac{d}{T}}$ which yields
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim\frac{1}{2}\sqrt{dT}$
which concludes our proof. ∎
@@
Furthermore, we assume that there exists some $d$ such that $\mathcal{M}_{\rho}(\epsilon,\Pi)\gtrsim\epsilon^{-d}$. Additionally, we know from Lemma B.1 that $\log\mathcal{N}_{2}(\epsilon,\Pi)\lesssim\epsilon^{d}$. We now provide a regret lower bound for Lipschitz bandits.
Lemma C.3.
The regret lower bound for the Lipschitz bandit is $\Omega\left(T^{\frac{d+1}{d+2}}\right)$.
@@
Due to Lemma B.1, the covering and packing number can now be bounded by
$\log\mathcal{N}_{2}\left(\sqrt{\frac{2}{\delta}},\Pi\right)$ $\lesssim\left(\frac{2}{\delta}\right)^{d/2}$
$\log\mathcal{M}_{\rho}(\epsilon,\Pi)$ $\gtrsim{\epsilon}^{-d}.$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}$ $\gtrsim 1-\left({1+\left(\frac{2}{\delta}\right)^{d/2}+T\delta}\right)\epsilon^{d}.$
We choose $\delta=2^{\frac{d}{d+2}}T^{-\frac{2}{d+2}}$. This yields that
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\left(1+(2T)^{\frac{d}{d+2}}\right)\epsilon^{d}.$
Selecting $\epsilon=\left(2T\right)^{-\frac{1}{d+2}}4^{-\frac{1}{d}}$ yields
$\mathcal{BR}^{*}(T)$ $\gtrsim 2^{-\frac{1}{d+2}}4^{-\frac{1}{d}}T^{\frac{d+1}{d+2}}-o(1)$
which yields the expected (and optimal) bound. ∎

### v3

@@ 7 Conclusions and Future Work
We introduce a general setting that embeds contextual information in a Bayesian setting. This setting covers a wide variety of online decision-making tasks. Using information-theoretic tools, we demonstrated a general method for obtaining regret lower bounds for problems in this setting. We used this method to present regret lower bound which depends on the information accumulated by the agent. We also presented regret upper bounds for Thompson sampling which depends on the accumulated information. These results quantify the relationship between a-priori external information and regret in online settings and the relationship between online information accumulation and regret. We then utilized this trade-off in a multiple-choice question-answering task with LLMs, demonstrating that the aforementioned quantification can be easily used in an online setting.
Limitations.
@@
A central assumption in our analysis is that exogenous information can be gathered regardless of the process history. However, this assumption may be violated in a general sequential decision process where the action itself may have ramifications on the quality, quantity, and cost of the exogenous information. Our work only covers the Bayesian setting, and our definition of the measure of information relies on it. In other settings, such as adversarial learning [Neu and Olkhovskaya, 2020] or frequentist settings, a different measure of information is required. Another limitation is the difference in $\sqrt{\log K}$ between the lower and upper bounds presented in this work. Making these bounds tighter can be the topic of future work.
@@
Finally, and importantly, with the increasing prevalence of LLMs, and foundation models in general, building solid foundations as well as practical algorithms for using prior knowledge in sequential decision-making is an important research endeavor that may be built upon the foundations that are laid in this paper.
@@ Appendix A Table of Notations
$R_{M}(\pi,C)$ reward distribution according to decision $\pi$ on model $M$.
$\mu_{M}(\pi,C)$ mean reward according to decision $\pi$ on model $M$.
@@
$\mathcal{BR}_{P}(T)$ The Bayesian regret given prior $P$.
$\mathcal{BR}^{*}(T)$ The worst-case Bayesian regret.
$\mathcal{N}(d,\epsilon,H)$ $\epsilon$-covering number of $H$ concerning metric $d$.
$\mathcal{M}(d,\epsilon,H)$ $\epsilon$-packing number of $H$ concerning metric $d$.
$\mathcal{M}^{\mathrm{loc}}(d,\epsilon,H)$ $\epsilon$-local packing number of $H$ concerning metric $d$.
$\mathrm{Vol}(\cdot)$ Volume of a set.
@@ Appendix B Useful Properties of Covering Numbers
Definition B.1 (Covering and packing numbers [Wainwright, 2019]).
@@
  • A set $C\subset H$ is said to be an $\epsilon$-covering set of $H$ if for all $h\in H$ there is $c\in C$ such that $d(h,c)\leq\epsilon$.
@@
    $\mathcal{M}(d,\epsilon,H)=\min\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is an $\epsilon$-covering set of $H$}\right\}.$
@@
  • A set $C\subset H$ is said to be a packing set of $H$ if for all $c_{1},c_{2}\in C$ we have that $d(c_{1},c_{2})\geq\epsilon$.
@@
    $\mathcal{N}(d,\epsilon,H)=\max\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is an $\epsilon$-packing set of $H$}\right\}.$
@@
  • We now define $B(\theta,\epsilon)=\{\theta^{\prime}\in H:d(\theta^{\prime},\theta)\leq\varepsilon\}$. $C\subseteq B$ is an $\epsilon$-local packing set of $B(\theta,\epsilon)$ if for all $c_{1},c_{2}\in C$, $d(c_{1},c_{2})\geq\frac{\epsilon}{2}$.
@@
    $\mathcal{M}^{\mathrm{loc}}(d,\epsilon,H)=\max\left\{m:\exists B(\theta,\varepsilon),\text{such that $C$ is an }\epsilon-\text{local packing of $B$ and }\left|C\right|=m\right\}.$
@@
  • (Theorem 14.2 of Wu [2020]) For any subset of dimension $d$ that is contained in $\mathbb{B}_{p}^{d}$ the covering number $\mathcal{N}_{p}(\epsilon)$ obeys $d\log\left(\frac{2}{\epsilon}\right)\leq\log\mathcal{N}_{p}(\epsilon)\leq d\log\left(\frac{3}{\epsilon}\right)$. This includes the probability simplex with $d$ variables and $\mathbb{B}_{d}^{2}$.
@@
    $\mathcal{M}(2\epsilon)\leq\mathcal{N}(\epsilon)\leq\mathcal{M}(\epsilon)$
Throughout the appendix, $\mathrm{Vol}(A)$ denotes the volume of set $A$.
@@ Appendix C Proofs for Section 3
$\mathcal{BR}^{*}(T)$ $=\sup_{\nu}\inf_{p_{1},\dots,p_{T}}\mathbb{E}_{M\sim\nu}\left[\sum_{t=1}^{T}(\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t},C_{t}))\right]$
$\geq\sup_{\nu}\inf_{p_{1},\dots,p_{T}}\frac{\epsilon T}{2}\mathbb{P}_{\nu}\left(\sum_{t=1}^{T}(\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t},C_{t}))\geq\frac{\epsilon T}{2}\right)$
$\geq\frac{\epsilon T}{2}\sup_{\nu}\inf_{p_{1},\dots,p_{T}}\mathbb{P}_{\nu}\left(\forall t:\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t},C_{t})\geq\frac{\epsilon}{2}\right)$
$\geq\frac{\epsilon T}{2}\max_{k\in\{1,\dots,K\}}\inf_{p_{1},\dots,p_{T}}\mathbb{P}\left(\forall t:\left|\mu_{M}(\phi_{k},C_{t})-\mu_{M}(p_{t},C_{t})\right|\geq\frac{\epsilon}{2}\right)$
$\geq\frac{\epsilon T}{2}\max_{k\in\{1,\dots,K\}}\inf_{p_{1},\dots,p_{T}}\mathbb{P}\left(\forall t:\rho(\phi_{k},\pi_{t})\geq\frac{\epsilon}{2}\right)$
$\geq\frac{\epsilon T}{2}\max_{k}\inf_{t}\mathbb{P}(p_{t}\neq\phi_{k})$
$\geq\frac{\epsilon T}{2}\left(1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}\right).$
The first inequality follows from Markov’s inequality, the second inequality from a union bound, and the last inequality from Fano’s inequality (Theorem 3.1). ∎
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}\right].$ (12)
    Applying Theorem 3.3 we have that $I(V;\mathcal{H}_{T})\leq 2\bar{A}\epsilon^{2}T$. Substituting this back into Equation 12,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+2\bar{A}\epsilon^{2}T}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}\right].$
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}\right].$ (13)
    Applying Theorem 3.3 we have that $I(V;\mathcal{H}_{T})\leq 2\bar{A}\epsilon^{2}T$. Substituting this back into Equation 13,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+2\bar{A}\epsilon^{2}T}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}\right].$
@@
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}(\rho,\Pi,\epsilon)}\right].$ (14)
    Applying Theorem 3.3 we have that $I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}\left(\rho,\Pi,\sqrt{\delta}\right)+T\delta\right)$. Substituting this back into Equation 14,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+\inf_{\delta>0}\left(\log\mathcal{N}\left(\rho,\Pi,\sqrt{\delta}\right)+T\delta\right)}{\log\mathcal{M}(\rho,\Pi,\epsilon)}\right].$
@@ C.3 Proofs for Additional Lower Bounds
Proposition C.1.
For any algorithm, there exists a Bayesian decision-making task with a finite decision set and a single context ($\left|\mathcal{C}=1\right|$) such that the regret can be lower bounded by $\Omega(\sqrt{\left|\Pi\right|T})$.
@@
We focus on the set $\Delta(\Pi)$ and note that it is a parametric set. We also see that for all policies $\phi,\psi\in\Delta(\Pi)$ we have that
$\mathbb{E}[\left|\mu_{M}(\phi,C)-\mu_{M}(\psi,C)\right|]=c\left\|\phi-\psi\right\|_{1}.$
for some positive constant $c$. Hence, in this scenario, we can take $\rho(x,y)=c\left\|x-y\right\|_{1}$. It is convenient to denote $K=\left|\Pi\right|$. Thus, using a scaling argument,
$\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$ $=\log\mathcal{M}^{\mathrm{loc}}_{1}(\Delta(\Pi),\epsilon/c).$
From the definition of a local packing number, we have that
$\log\mathcal{M}^{\mathrm{loc}}_{1}(\Delta(\Pi),\epsilon/c)$ $\geq\log\left(\frac{\mathrm{Vol}(\mathbb{B}_{1}^{K-1}(\epsilon))}{\mathrm{Vol}(\mathbb{B}_{1}^{K-1}(\epsilon/2))}\right)$
$\geq(K-1)\log\frac{2\epsilon}{\epsilon}$
$=K-1.$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+2\bar{A}\epsilon^{2}T}{K-1}\right].$
Now, we select $\epsilon$ to maximize such that our lower bound will be the tightest. In particular, we choose $\epsilon=\sqrt{\frac{K-1}{6T\bar{A}}}$ which yields the following lower bound,
$\mathcal{BR}^{*}(T)$ $\geq\frac{1}{2}\sqrt{\frac{T(K-1)}{6\bar{A}}}\left[\frac{2}{3}-\frac{1}{K-1}\right].$
This concludes our proof. ∎
The results for multi-armed bandits and tabular MDPs are shown by substituting $\left|\Pi\right|=K$ and $HSK$ respectively.
Proposition C.2.
For any algorithm, there exists a Bayesian decision-making task with $\Pi\subseteq\mathbb{B}_{2}^{d}$, where the reward means are a linear function of the decisions, such that the regret can be lower bounded by $\Omega(\sqrt{dT})$.
@@
Similarly to Proposition C.1, we have $\rho(x,y)=c\left\|x-y\right\|_{1}$ for some positive constant $c>0$. We focus on the set $\mathbb{B}_{2}^{d}$ and note that it is a parametric set. Furthermore, using Lemma B.1 we have that
$\log\mathcal{M}^{\mathrm{loc}}(\rho,\mathbb{B}_{2}^{d},\epsilon)\geq d.$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\frac{1+2\bar{A}\epsilon^{2}T}{d}\right].$
Now, we select $\epsilon$ to maximize such that our lower bound will be the tightest. In particular, we choose $\epsilon=\sqrt{\frac{d}{6T\bar{A}}}$ which yields the following lower bound,
$\mathcal{BR}^{*}(T)$ $\geq\frac{1}{2}\sqrt{\frac{Td}{6\bar{A}}}\left[\frac{2}{3}-\frac{1}{d}\right].$
This concludes our proof. ∎
@@
Unlike the previous settings, we note that the decision set is not parametric this time. We now prove a lower bound for this setting.
Proposition C.3.
The regret lower bound for the Lipschitz bandit is $\Omega\left(T^{\frac{d+1}{d+2}}\right)$, where $d$ is chosen such that
$c_{1}\leq\log\mathcal{N}(\rho,\epsilon,\Pi)\cdot\epsilon^{d}\leq c_{2}$
@@
From the assumption, the covering and packing number can now be bounded by
$\log\mathcal{N}\left(\rho,\sqrt{\frac{2}{\delta}},\Pi\right)$ $\leq c_{2}\left(\frac{2}{\delta}\right)^{d/2}$
$\log\mathcal{M}(\rho,\epsilon,\Pi)$ $\geq c_{1}{\epsilon}^{-d}.$
@@
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}$ $\geq\frac{1}{2}\left[1-\left({1+c_{2}\left(\frac{2}{\delta}\right)^{d/2}+T\delta}\right)\frac{\epsilon^{d}}{c_{1}}\right].$
We choose $\delta=T^{\frac{2}{d+2}}c_{2}^{-\frac{2}{d+2}}2^{-\frac{d}{d+2}}$. This yields that
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq\frac{1}{2}\left[1-\left(1+2(2c_{2})^{-\frac{d}{d+2}}T^{\frac{d}{d+2}}\right)\frac{\epsilon^{d}}{c_{1}}\right].$
Selecting $\epsilon^{d}=\frac{1}{2}\left[\frac{c_{1}}{\left(1+(2c_{2})^{-\frac{d}{d+2}}T^{\frac{d}{d+2}}\right)}\right]$ yields
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}$ $\geq\frac{1}{4}$
Now, $\epsilon T\gtrsim T^{\frac{d+1}{d+2}}$ so we have that
$\mathcal{BR}^{*}(T)\gtrsim T^{\frac{d+1}{d+2}}.$
@@
C.4 Additional Explanation
We now provide an additional explanation regarding Proposition 3.4. In particular, we explain why we had to select a local packing set for parametric decision spaces $\Pi$ and why we focused on the simplex of the decision set for finite decision spaces.
Selecting $\Delta(\Pi)$ over $\Pi$.
This explanation is rather straightforward. We see that by selecting $\Pi$ we have that $\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)=1$ for any value of $\epsilon$ which makes the lower bound null.
Selecting a local-packing set for parametric $\Pi$.
This decision stems from the improved upper bound of $I(V;\mathcal{H}_{T})$ for parametric decision space, which is found in Theorem 3.3. Using the other upper bound for $I(V;\mathcal{H}_{T})$ simply results in a sub-optimal lower bound. Additional details can be found in Yang and Barron [1999].